OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 4 cm and a height of 12 cm. The dimensions of the actual water tanks that will be installed in the park are 5 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2. Use 22 over 7 for pi.

OpenStudy (anonymous):

you have surface area of cylinder as \[2\pi r ^{2}(r+h)\] and you have dimensions as R=5*r and H=5*h solve it by your self

OpenStudy (anonymous):

my answer is 7536?????????????????????

OpenStudy (anonymous):

the dimesions of large tank is 5 times smaller one its height=60 cm radius=20cm ttotal surface area=2*pie*r*h+pie*r*r* 2*22/7*20*60+22/7*20*20 =8792

OpenStudy (anonymous):

oops i put wrong formula sorry guys. . . \[2 \pi r(r+h)\] let him solve by himself just tell the formula. . .

OpenStudy (anonymous):

@theyatin wtf I don't undertstand explain

OpenStudy (anonymous):

ok 2πr(r+h) is the formula of surface area of cylinder and its is given that original cylender has five times as dimensions of model so R=5r(5*4cm) and H=5h(5*12cm) hence put values and solve

OpenStudy (anonymous):

all I see is number and letters

OpenStudy (anonymous):

ohk 2πr(r+h)= 2* 22/7* 20cm (20cm+60 cm)= solve it